Hi there. Can "x^2 - 10x + 22 = -2" be solved by taking the square root? Can "4x^2 + 8x + 3 = 0" be solved by taking the square root? If so, please show me an explanation.
Thank you for all your help! :D

Answer by TimothyLamb(4379)

x^2 - 10x + 22 = -2
x^2 - 10x + 24 = 0
---
the above quadratic equation is in standard form, with a=1, b=-10 and c=24
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -10 24
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
answer:
x = 6
x = 4
---
4x^2 + 8x + 3 = 0
---
the above quadratic equation is in standard form, with a=4, b=8 and c=3
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
4 8 3
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
answer:
x = -0.5
x = -1.5
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---